140817 LML Observations

Isaac arrived at LML at 9:30am. The doors to the collector had been deployed but no sample was present in the collection jar. The collector was then disassembled and the body of the collector as well as its components were sprayed initially with 409 followed by a rinse with DI water. The collector was then reassembled for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *